OpenStudy (moongazer):

how to find the general solution of this: cos y sin 2x dx + ( (cos y)^2 -(cos x )^2)dy = 0 I am trying substitution but I can't find anything that will lead me to a solution.

OpenStudy (moongazer):

anyone?

OpenStudy (moongazer):

yes

OpenStudy (moongazer):

the answer is: [cos x(1 + siny)]^2 = cos y(y+c-cos y) I don't know how to get there

OpenStudy (moongazer):

sorry but I can't understand your solution could you explain it further? thanks :)

OpenStudy (moongazer):

isn't it that you can't integrate it immediately ?

OpenStudy (anonymous):

\[\cos y\sin2x~dx+\left(\cos^2y-\cos^2x\right)~dy=0\] Check to see if it's exact: \[\begin{cases}M(x,y)=\cos y\sin2x\\N(x,y)=\cos^2y-\cos^2x\end{cases}~~\Rightarrow~~\begin{cases}\dfrac{\partial M}{\partial y}=-\sin y\sin 2x\\\\ \dfrac{\partial N}{\partial x}=2\cos x\sin x=\sin2x \end{cases}\] the equation isn't exact, but we can find an integrating factor in terms of \(y\): \[\begin{align*}\ln\mu(y)&=\int\frac{\dfrac{\partial N}{\partial x}-\dfrac{\partial M}{\partial y}}{M}~dy\\\\ &=\int\frac{\sin2x+\sin y\sin2x}{\cos y\sin2x}~dy\\\\ &=\int\frac{1+\sin y}{\cos y}~dy\\\\ &=\int(\sec y+\tan y)~dy\\\\ &=\ln|\sec y+\tan y|-\ln|\cos y|\\\\ &=\ln\left|\frac{\sec y+\tan y}{\cos y}\right|\\\\ \mu(y)&=\frac{\sec y+\tan y}{\cos y}\\\\ &=\frac{1+\sin y}{\cos^2y} \end{align*}\]Now you should have an exact equation: \[\begin{align*}\frac{1+\sin y}{\cos^2y}\cos y\sin2x~dx+\frac{1+\sin y}{\cos^2y}\left(\cos^2y-\cos^2x\right)~dy&=0\\\\ \frac{\sin2x+\sin2x\sin y}{\cos y}~dx+\left(1+\sin y-\frac{1+\sin y}{\cos^2y}\cos^2x\right)~dy&=0 \end{align*}\] So now, \[\begin{cases}M=\dfrac{\sin2x+\sin2x\sin y}{\cos y}\\\\N=1+\sin y-\dfrac{1+\sin y}{\cos^2y}\cos^2x\end{cases}~~\Rightarrow~~\begin{cases}\dfrac{\partial M}{\partial y}=\dfrac{\sin2x(1+\sin y)}{\cos^2y}\\\\ \dfrac{\partial N}{\partial x}=\dfrac{\sin2x(1+\sin y)}{\cos^2y} \end{cases}\] which means the equation is exact.

OpenStudy (anonymous):

We're looking for a solution of the form \(\Psi(x,y)=C\), where \(\dfrac{\partial \Psi}{\partial x}=M(x,y)\) and \(\dfrac{\partial \Psi}{\partial y}=N(x,y)\). \[\begin{align*} \dfrac{\partial\Psi}{\partial x}&=M\\\\ \partial\Psi&=\frac{(1+\sin y)\sin2x}{\cos y}~\partial x\\\\ \int\partial\Psi&=\frac{1+\sin y}{\cos y}\int\sin2x~\partial x\\\\ \color{red}\Psi&\color{red}{=-\frac{(1+\sin y)\cos2x}{2\cos y}+f(y)}\\\\ \frac{\partial\Psi}{\partial y}&=-\frac{1}{2}\frac{\partial}{\partial y}\left[\frac{(1+\sin y)\cos2x}{\cos y}\right]+f'(y)\\\\ N(x,y)&=-\frac{\cos2x(1+\sin y)}{2\cos^2y}+f'(y)\\\\ 1+\sin y-\dfrac{1+\sin y}{\cos^2y}\cos^2x&=-\frac{\cos2x(1+\sin y)}{2\cos^2y}+f'(y)\\\\ f'(y)&=(1+\sin y)\left(1-\frac{\cos^2x}{\cos^2y}+\frac{\cos2x}{2\cos^2y}\right)\\\\ f'(y)&=(1+\sin y)\left(1-\frac{1+\cos2x}{2\cos^2y}+\frac{\cos2x}{2\cos^2y}\right)\\\\ f'(y)&=(1+\sin y)\left(1-\frac{1}{2\cos^2y}\right)\\\\ f'(y)&=1+\sin y-\frac{1}{2}\sec^2y-\frac{1}{2}\frac{\sin y}{\cos^2y}\\\\ f(y)&=\int\left(1+\sin y-\frac{1}{2}\sec^2y-\frac{1}{2}\sec y\tan y\right)~dy\\\\ \color{blue}{f(y)}&\color{blue}{=y-\cos y-\frac{1}{2}\tan y-\frac{1}{2}\sec y+C} \end{align*}\] So the final solution should be \[\Psi(x,y)=-\frac{(1+\sin y)\cos2x}{2\cos y}+y-\cos y-\frac{1}{2}\tan y-\frac{1}{2}\sec y+C=0\] or simply \[-\frac{(1+\sin y)\cos2x}{2\cos y}+y-\cos y-\frac{1}{2}\tan y-\frac{1}{2}\sec y=C\]